Find the equilibrium price and quantity
Set Qd = Qs:
100/3-P/3 = -5+P/2
1
Expert's answer
2017-12-10T07:07:06-0500
100/3-P/3 = -5+P/2
100/3+5=P/3+P/2
200+30=2P+3P
230=5P, => P=46, Q=18